# Augmentation-based Unsupervised Cross-Domain Functional MRI Adaptation for Major Depressive Disorder Identification

### v1

@@ VI-A Ablation Experiment
VI-B Impact of Hyperparameters
[…]
@@ VI-B Impact of Hyperparameters
VI-C Discriminative Brain Regions
This work addresses the problem of data heterogeneity in multi-site fMRI while identifying the most discriminative brain regions in the task of MDD vs. NC classification. In Section IV-B1, the self-attention mechanism we employed to learn the attention matrix $A$ can efficiently quantify the interdependence of different brain regions. Given that our proposed AUFA framework is used for brain disorder detection tasks, the elements with larger values in $A$ indicate those brain region connections that play a decisive role in distinguishing healthy from diseased states. To gain deeper insight into these critical connections, we averaged the multiple attentional matrices learned for each attention head, aiming to strengthen those significant connections consistently apparent across multiple attentional perspectives. Subsequently, the top ten most discriminative brain region connections were visualized by the BrainNet Viewer [51] tool, and the results are shown in Fig. 5. As shown in Fig. 5, the identified brain regions such as the calcarine fissure (CAL), superior temporal gyrus (STG), lingual gyrus (LING), and medial and cingulum gyrus (DCG) play important roles. This is highly correlated with previous studies of MDD [52, 53, 54, 55], thus validating the effectiveness of AUFA based on rs-fMRI in identifying disease-related functional connectivity patterns. In the Supplementary Materials, we also elaborate on the influence of different source domains.
@@ VI-D Limitations and Future Work
The current work still has several limitations. First, The proposed framework only considers resting-state fMRI data, while other modalities (e.g., structural MRI and demographic information) can provide rich complementary information about the brain from different perspectives. Employing multi-modality data to help boost learning performance, which will also be our further work. Second, this work mainly focuses on dealing with the migration problem from a single source domain to a single target domain. We can develop more advanced models to integrate and optimize information from different source domains, and explore effective adaptive strategies for complex scenarios involving multiple source domains and multiple target domains. In addition, it is interesting to generalize the model to unknown domains. In the future, we will focus on constructing and optimizing generative models for modeling data differences between source and target domains to facilitate knowledge migration and generalization, and to further improve the generalization performance of our method over unknown domains.
Figure 5: Visualization of the top ten most discriminative functional connections detected by our method in the task of MDD vs. NC classification using the BrainNet Viewer [51].
@@ VII Conclusion
In this paper, we propose an augmentation-based unsupervised cross-domain fMRI adaptive framework for the automatic diagnosis of major depressive disorder (MDD). We aim to reduce the distributional differences of fMRI data from different sites and address the overfitting of the model in the source domain. We learn graph-level features with spatial attention via Transformer encoder and mitigate data heterogeneity between domains via the domain adaptation module. We also augment the target domain data to further improve the generalization of the model in the target domain. Experimental results on multiple sites demonstrate the effectiveness of AUFA in MDD detection. In addition, our framework can discover discriminative brain regions, providing potential imaging biomarkers for MDD analysis.

### v2

@@ VI-B Impact of Hyperparameters
VI-C Discriminative Brain Regions
This work addresses the problem of data heterogeneity in multi-site fMRI while identifying the most discriminative brain regions in the MDD vs. NC classification task. In Section IV-B1, we utilize self-attentive mechanisms to learn that the attention matrix $A$ can effectively quantify the interdependence between different brain regions. Given that our proposed AUFA framework is used for brain disorder detection tasks, the elements with larger values in $A$ indicate those brain region connections that play a decisive role in distinguishing healthy from diseased states. To gain deeper insight into these critical connections, we averaged the multiple attentional matrices learned for each attention head, aiming to strengthen those significant connections consistently apparent across multiple attentional perspectives. Subsequently, the top ten most discriminative brain region connections were visualized by the BrainNet Viewer [51] tool, and the results are presented in Fig. 5. As can be seen in Fig. 5, the identified brain regions such as the calcarine fissure (CAL), superior temporal gyrus (STG), lingual gyrus (LING), and medial and cingulum gyrus (DCG) play important roles. This is highly correlated with previous studies of MDD [52, 53, 54, 55], thus validating the effectiveness of AUFA based on rs-fMRI in identifying disease-related functional connectivity patterns. In the Supplementary Materials, we also elaborate on the influence of different source domains.
@@ VI-C Discriminative Brain Regions
Figure 5: Visualization of the top ten most discriminative functional connections detected by AUFA in the MDD vs. NC classification task using the BrainNet Viewer [51].
@@ VI-D Limitations and Future Work
There are still several limitations to the present study. First, The proposed framework only considers resting-state fMRI data, while there exists untapped potential in integrating other modalities like structural MRI and demographic information. These additional sources offer valuable, diverse insights into the brain, enriching our understanding from multiple viewpoints. We will employ multi-modal data in the future to help boost learning performance. Second, this work mainly focuses on dealing with the migration problem of a single source/target domain. We can develop more advanced models to integrate and optimize information from multiple source domains, and explore effective adaptive strategies for complex scenarios involving multiple source/target domains. In addition, it is interesting to generalize the model to unknown domains. In the future, we will focus on constructing and optimizing generative models for modeling data differences between source and target domains. This will facilitate the migration of knowledge, thereby enhancing the overall generalization performance of our approach across unknown domains.
@@ VII Conclusion
In this paper, we construct an augmentation-based unsupervised cross-domain fMRI adaptive framework for the automatic diagnosis of major depressive disorder (MDD). We aim to reduce the distributional differences of fMRI data from different sites and address the overfitting of the model in the source domain. We learn graph-level features with spatial attention via Transformer encoder and mitigate data heterogeneity between domains via the domain adaptation module. We also augment the target domain data to enhance the model’s generalization in that specific domain. Experimental results on multiple sites validate the effectiveness of AUFA in MDD detection. In addition, our framework can discover discriminative brain regions, providing potential imaging biomarkers for MDD analysis.